OpenStudy (anonymous):

Combine like terms. 5/6x - 2/3x A. 1/5x B. 1/3x C. 1/2x D. x

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

what is \[\frac{5}{6}-\frac{2}{3}\]?

OpenStudy (anonymous):

hold up a sec XD

OpenStudy (anonymous):

1 1/2?

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):

common denominator is 6

OpenStudy (anonymous):

opps my bad, i added XD

OpenStudy (anonymous):

its 1/6

OpenStudy (misty1212):

ok try again

OpenStudy (misty1212):

ok good, now stick an \(x\) next to it

OpenStudy (misty1212):

i see that is not one of your answers, so i hope there is a typo there

OpenStudy (anonymous):

ok but none of my answers fit with this one

OpenStudy (misty1212):

answer should be \[\frac{5}{6}x-\frac{2}{3}x=\frac{1}{6}x\]

OpenStudy (anonymous):

opps it was a typo XD

OpenStudy (anonymous):

you were correct XD

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]